Hellllpppp
A wire Is 1.1 m long and 1.1 mm^2 in cross-section
Answer #1

Given wire lenght L = 1.1 m, cross-sectional area A = 1.1*10^-6 m2, current I = 5.6 A, potential difference V = 2.1 V

we know that from Ohm's law V = IR ==> R = V/I = 2.1/5.6 =0.375 ohms
and resistivity rho = R*A/L = 0.375*1.1*10^-6 /1.1 = 375000 ohm m

now conductivity sigma = 1/ rho = 1/375000 = 2.67*10^-6 (ohm m )^-1